If A+B+C=22/7Prove that -. Cos^2 a/2 + cos^2 b/2+ cos^2 c/2=2 (1+ sin a/2. + Sin b/2 + sin c/2 )

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Lhs = 3*(3/4 )
Keeping all angles 60 like for an equilateral triangle

Rhs = 2*(1+ 3/2) = 5

Kindly check and repost correct question

Regards